a,b,c,d are all real numbers and if (a2+b2+c2)p2 -2(ab+bc+cd)p + b2+c2+d2 <= 0 then prove that a,b,c,d are in GP


 


 

Grade:12

1 Answers

jitender lakhanpal
10 years ago

hi rohit

the discriminant of the equation cant be -ve as the leading coeff is +ve so its an upward parabola

which cuts at 2 points in x axis

now finding the discriminant

(ab+bc+cd)2   - (a2+b2+c2 )  (b2+c2+d2)     > = 0

but the L.H.S is not greater than 0 because of CAUCHY SCHWARZ INEQUALITY

so it can only be 0

so ∑ a2  ∑ b2   = ∑ ab                   applying for a , b ,c  and b , c, d

and in equality is true only when the terms involved in the sigma sign are proportional

i.e   a/b = b/c = c/d     so proved they are in G.P

We are all IITians and here to help you in your IIT JEE preparation.

Now you can win by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar.. 

 

Thanks and Regards

Jitender

(askIITians expert)